OpenStudy (anonymous):

let f(x)= sin (2x) give the value of lim as h approaches 0; f(h) - f (0)/h

OpenStudy (iouri.gordon):

are you asking for \[\lim_{h \rightarrow 0}(\sin2h-\sin0)/h\] ?

OpenStudy (anonymous):

We Know the fact that \[\lim_{x \rightarrow 0}\frac{ \sin x }{ x } = 1\] to use this equation, transform formulus. \[\lim_{h \rightarrow 0}\frac{ \sin 2h }{2h } * 2 \] now, you result in solution.